OpenStudy (anonymous):

Damien works part-time from home as a data entry professional. He enters data in forms and is paid based on the entry accuracy rate. The table below shows his compensation. accuracy Comp per form tier1 14$ tier 2 12$ tier 3 0$ Based on his regular performance, the equation below can be used to determine C, the total amount Damien earns a month entering data, where x represents the total number of forms he completes that month. C = 14(0.6x) + 12(0.3x) + 0(0.1x) What is the minimum number of forms he must complete in a month to earn a total of at least $1,080?

OpenStudy (anonymous):

these are the only four answers it might be 91 90 1092 1068

OpenStudy (lilg132):

1,080 = 14(0.6x) + 12(0.3x) + 0(0.1x) 1080 = 8.4x + 3.6x + 0 1080 = 12x 1080 / 12 = 12x / 12 x = 90 he must complete 90 forms :)

OpenStudy (anonymous):

Thanks. Sorry I got like 5 more uhhh

OpenStudy (lilg132):

no problem i dont mind helping you

OpenStudy (anonymous):

thanks you dont know how much